How many liters of pure antifreeze (100% solution) must be added to 30 liters of a 15% antifreeze solution to get a 40% antifreeze solution?
Answer by stanbon(75887)

How many liters of pure antifreeze (100% solution) must be added to 30 liters of a 15% antifreeze solution to get a 40% antifreeze solution?
---
alcohol + alcohol = alcohol
----
1*x + 0.15*30 = 0.40(x+30)
----
Multiply thru by 100 to get:
100x + 15*30 = 40*x + 40*30
60x = 25*30
x = (1/2)25 = 12.5 liters (amt. of pure antifreeze needed)
============================================================
cheers,
Stan H.